A loan charging 6% interest compounded semi-annually, was repaid in three installments in 1 year: one $300 payment after 6 months; one $500 payment after 8 months; and a final payment of $1200 at the end of the year. What was the value of the original loan?
Answer #1
Interest rate for a month = 6%/12 m = 0.5% per month i.e. 0.005
Period Cash Flow PV Factor @0.5% Present Value
a b c d=b*c
First 6 month $             300 0.9705181 $         291.155
First 8 month $             500 0.9608852 $         480.443
First 12 month $         1,200 0.9419053 $      1,130.286
Total $      1,901.884
Loan Amount = $1901.8844
Loan Amount = $1901.88 (rounded off )
Working Note:
PV Factor
First 6 month =1/1.005^6 0.970518
First 8 month =1/1.005^8 0.960885
First 12 month =1/1.005^12 0.941905